If your grandparents deposited $5000 in a certificate on the day you were born at a rate of 4.5%, how much money would you have to start college after your 18th birthday if it was compounded monthly?
Answer by stanbon(57346)

If your grandparents deposited $5000 in a certificate on the day you were born at a rate of 4.5%, how much money would you have to start college after your 18th birthday if it was compounded monthly?
------------------------------------
A(t) = P(1+(r/n))^(nt)
A(18) = 5000(1+(0.045/12))^(12*18)
A(18) = 5000*2.24451
A(18) = $11222.53
----------------
Cheers,
Stan H.